Let X be a r.v. with probability density function f(x)-e(4-x2), -2 < otherwise (a) What is the value of c? (b) What is the cumulative distribution function of X? (c) What is EX) and VarX
Answer #1

a)

e(4-2) dr= 33-1

c = 3/32

b)

3r 31 F(r)=/3 (4 - 2) dr -8-32+2

c)

2 32

32 E(X2)-1%(4-2) dr-4 За (Decimal 0.8) 2 32

Var(X) = E(X^2)- E(X)^2 = 0.8

Please rate